Set up the equation for the following situation and solve.
Adam has three times as much money as Betty. Together they have $34. How much money does each person have ?

Answer by DrBeeee(684)

Let A = amount of money Adam has.
Let B = amount of money Betty has.
Then you have two equations
(1) A = 3*B and
(2) A + B = 34
Now substitute (1) into (2) and get
(3) 3*B + B = 34 or
(4) 4*B = 34 or
(5) B = 8.5
and from (1) we get
(6) A = 3*8.5 or
(7) A = 25.5
Let's check this with (2).
Is (25.5 + 8.5 = 34)?
Is (34 = 34)? Yes
Answer: Adam has $25.50 and Betty has $8.50